To make modern electronics components, silicon of extremely high purity is required. One of the steps in purifying silicon involves reacting the intermediate silicon tetrachloride with steam to produce high-purity silicon dioxide and hydrogen chloride gas, as shown in the unbalanced equation below.

SiCl4(s) + H2O(g) → SiO2(s) + HCl(g)

If the reaction is run with 410.7 g SiCl4 and 105.1 g H2O, what mass of SiO2 will be produced, and what is the excess reactant and what mass of it remains?

(a) What mass of SiO2 will be produced?
(b) What is the excess reactant? (Omit states-of-matter from your answer.)

(c) What mass of the excess reactant remains?
Answer #1

The balanced equation is SiCl4(s) + 2H2O(g) → SiO2(s) + 4HCl(g)

Molar mass (g/mole)         167          18           60              36.5

From the balanced reaction,

1 mole = 167 g of SiCl4 reacts with 2 moles = 2*18g = 36 g of H2O

410.7 g of SiCl4 reacts with (410.7g*36g)/167g = 88.5g of H2O

So 105.1g - 88.5g = 16.6 g of H2O left unreacted so it is the excess reactant.

(b) Therefore the excess reactant is H2O

(c) Therefore the mass of excess reactant H2O remains is 16.6g

(a) 1 mole = 167 g of SiCl4 produces 1 mole = 60 g of SiO2

410.7 g of SiCl4 produces (410.7g*60g)/167g = 147.6g of SiO2

Therefore the mass of SiO2 produced is 147.6 g